OpenStudy (anonymous):

Solve for b. a fraction showing five times b as the numerator and three as the denominator all equal to thirty

OpenStudy (imstuck):

\[\frac{ 5b }{ 3 }=30\]

OpenStudy (imstuck):

First thing to do is to get rid of the 3 in the denominator by multiplying it by both sides, like this:

OpenStudy (imstuck):

\[(\frac{ 3 }{ 1 })\frac{ 5b }{ 3 }=30(3)\]The 3's on the left side cancel out, leaving\[5b=90\]

OpenStudy (imstuck):

Divide both sides by 5 and you can solve for b. Do you know?

OpenStudy (imstuck):

90/5 = ?

OpenStudy (anonymous):

18

OpenStudy (imstuck):

Yep!!!!! That's great! Good job!

OpenStudy (anonymous):

yayy thanks

OpenStudy (imstuck):

TY for the medal!

OpenStudy (anonymous):

thank you